OpenStudy (anonymous):

Two cyclists drive a distance of 15 kilometres. One of the cyclists drives 10km h faster than the other one and arrives at their destination 20 minutes earlier. What are the speeds of the cyclists?

OpenStudy (lgbasallote):

x = speed of one 10 + x = speed of the other the one with 10 km/min +x speed got to the destination in 20 minutes so... 15kilometer/20 minutes = 10 km/hr + x let's convert km/min to km/hr first.. 15 km 60 min ---------- x --------- 20 minutes 1 hour 75 km/hr will be the converted value (see calculator) so... 75 km/hr -10 km/hr = x 65 km/hr = x x + 10 = 75 km/hr

OpenStudy (perl):

I used d = r*t let r be the speed of the slower cyclist, and t be the time for the slower cyclist. so we have 15 = r*t 15 = ( r + 10) ( t - 20/60) , since the units are km/hr , then 20 min = 20/60 hr

OpenStudy (perl):

so t = 15/r , and substitute that into second equation 15 = ( r + 10) ( 15/r - 1/3) , solve that